OpenStudy (anonymous):

Determine if 7, 8, and 13 can be the lengths of the sides of a triangle. If yes, classify the triangle

OpenStudy (anonymous):

You can use the Pythagorean Theorem to figure this out. The 2 shorter legs squared will equal the longest side (Hypotenuse) A^2 + b^ 2 = C^2 So 7^2 + 8^2 = 13^2 Is this true? If so, then your answer is yes.

OpenStudy (anonymous):

As for classifying the triangle, I am unsure.

OpenStudy (triciaal):

3 unequal lengths , 7, 8, 13 = scalene triangle other classifications 2 equal lengths = isosceles 1 angle more than 90° = obtuse 1 90° angle = right angle 3 equal sides/ 3 equal angles = equilateral